# The radius of statistical efficiency

### v1

@@ B.2 Proof of Theorem 2
In this section, we prove Theorem 2, or rather a stronger version thereof. To this end, we fix a $C^{1}$-smooth map $F\colon\mathbf{R}^{d}\to\bm{S}^{k}_{+}$ satisfying that it is integrable with respect to any measure $\mu\in\mathcal{P}_{2}$. We define the function $\mathcal{J}\colon\mathcal{P}_{2}\to\mathbf{R}$ by setting
@@
The differential of $F$ will be denoted by $DF(x)\colon\mathbf{R}^{d}\to\bm{S}^{k}$, while the symbol $DF(x)^{*}\colon\bm{S}^{k}\to\mathbf{R}^{d}$ will denote the adjoint linear map of $DF(x)$. We further assume that there exists a constant $L>0$ such that $\|DF(x)\|_{\rm op}\leq L(1+\|x\|)$ for all $x.$
In order to simplify notation, for any matrix $A\in\bm{S}^{k}$, we let $E_{k}(A)$ denote the set of all unit eigenvectors of $A$ corresponding to the minimal eigenvalue $\lambda_{\min}(A)$. We will also use the elementary fact that $\lambda_{\min}$ is a concave function on $\bm{S}^{k}$ and its supdifferential at any matrix $A$ is the set
$\partial\lambda_{\min}(A)=\operatorname{conv}\{uu^{\top}:u\in E_{k}(A)\}.$
See for example [9, Corollary 5.2.3, Corollary 5.2.4 (iii)]. Abusing notation, we will set $E_{k}(\mu):=E_{k}(\operatorname{\mathbb{E}}_{\mu}F(x))$ for any measure $\mu\in\mathcal{P}_{2}$.
The proof of Theorem 2 will be subdivided into two parts, corresponding to the two inequalities in (15). We begin by establishing the first inequality ${\rm REG}(\mu)^{q_{2}-1}\lesssim{\rm RSE}(\mu)$. The proof amounts to simply applying the fundamental theorem of calculus to the function $\mathcal{J}$ along a geodesic $\mu_{t}$ joining a measure $\mu$ to its nearest point in $\mathcal{E}$. This is the content of the following theorem.
@@
Fix a measure $\mu\in\mathcal{P}_{2}$, constants $c,\varepsilon>0$, and a power $q\in[0,1)$. Suppose that for all measures $\nu$ satisfying $W_{2}(\nu,[\mathcal{J}=0])\leq W_{2}(\mu,[\mathcal{J}=0])+\varepsilon$, the estimate holds:
$\min_{u\in E_{k}(\nu)}~{}\operatorname{\mathbb{E}}_{\nu}\|DF(x)^{*}[uu^{\top}]\|^{2}_{2}\leq c\cdot\mathcal{J}(\nu)^{2q}.$
Then, the inequality holds:
$W_{2}(\mu,[\mathcal{J}=0])\geq\tfrac{1}{(1-q)\sqrt{c}}\cdot\mathcal{J}(\mu)^{1-q}.$
@@
Fix a measure $\nu\in[\mathcal{J}=0]$ satisfying $W_{2}(\mu,\nu)\leq W_{2}(\mu,[\mathcal{J}=0])+\varepsilon$ and let $\pi\in\Pi(\nu,\mu)$ be an optimal transport plan between $\nu$ and $\mu$. Define the functions $\pi_{t}(x,y)=(1-t)x+ty$. Then the curve $\mu_{t}=(\pi_{t})_{\#}\mu$ is a constant speed geodesic between $\mu$ and $\nu$ [59, Theorem 5.27]:
$W_{2}(\mu_{t},\mu_{s})=|t-s|\cdot W_{2}(\mu,\nu)\qquad\forall t,s\in[0,1].$
Define the curve of matrices $\gamma(t)=\operatorname{\mathbb{E}}_{\mu_{t}}F(x)$. We would like to now compute $\dot{\gamma}(t)$ by exchanging differentiation and integration in the expression:
$\dot{\gamma}(t)=\frac{d}{dt}\mathop{\operatorname{\mathbb{E}}}_{(x,y)\sim\pi}F((1-t)x+ty).$ (75)
To this end, we bound the derivative of the integrand:
$\|DF((1-t)x+ty)[y-x]\|_{\rm op}\leq L(1+\|x\|_{2}+\|y\|_{2})\|y-x\|_{2}.$
Applying Hölder’s inequality, we see that the right side is $\pi$-integrable. Therefore, exchanging integration and differentiation in (75) yields the expression $\dot{\gamma}(t)=\mathop{\operatorname{\mathbb{E}}}_{(x,y)\sim\pi}DF((1-t)x+ty)[y-x]$.
It is straightforward to see that $\gamma$ is absolutely continuous and therefore using the subdifferential chain rule for concave functions [11, Lemma 3.3, p. 73], we deduce that for almost every $t\in(0,1)$ we have
@@
In particular, for each such $t$ we may choose $U_{t}$ satisfying the running assumption $\operatorname{\mathbb{E}}_{\mu_{t}}\|DF(x)^{*}[U_{t}]\|^{2}_{2}\leq c\cdot\mathcal{J}(\mu_{t})^{2q}$. Continuing with (76), we successively compute
$\frac{d}{dt}\mathcal{J}(\mu_{t})$ $=\mathop{\operatorname{\mathbb{E}}}_{(x,y)\sim\pi}\langle DF((1-t)x+ty)^{*}[U_{t}],y-x\rangle$ (77)
$\leq\mathop{\operatorname{\mathbb{E}}}_{(x,y)\sim\pi}\|DF((1-t)x+ty)^{*}[U_{t}]\|_{2}\cdot\|y-x\|_{2}$ (78)
$\leq\sqrt{\mathop{\operatorname{\mathbb{E}}}_{(x,y)\sim\pi}\|DF((1-t)x+ty)^{*}[U_{t}]\|^{2}_{2}}\cdot\sqrt{\mathop{\operatorname{\mathbb{E}}}_{(x,y)\sim\pi}\|y-x\|^{2}}$ (79)
$\leq\sqrt{c}\cdot W_{2}(\mu,\nu)\cdot\mathcal{J}(\mu_{t})^{q},$ (80)
where (79) follows from Hölder’s inequality. Raising $\mathcal{J}(\mu_{t})$ to power $1-q$, we deduce
$\frac{d}{dt}\mathcal{J}(\mu_{t})^{1-q}\leq\sqrt{c}(1-q)\cdot W_{2}(\mu,\nu)\qquad\textrm{for a.e. }t\in(0,1).$
Integrating both sides from $t=0$ to $t=1$, we conclude
$\mathcal{J}(\mu)^{1-q}-\mathcal{J}(\nu)^{1-q}\leq\sqrt{c}(1-q)\cdot W_{2}(\mu,\nu).$ (81)
Taking into account the equality $\mathcal{J}(\nu)=0$ and the estimate $W_{2}(\mu,\nu)\leq W_{2}(\mu,[\mathcal{J}=0])+\varepsilon$, we may now let $\varepsilon$ tend to zero 00 thereby completing the proof. ∎
@@
$|\nabla f|(x)=\limsup_{x^{\prime}\to x}\frac{(f(x)-f(x^{\prime}))_{+}}{\mathsf{d}(x,x^{\prime})}.$
Importantly, if a slope is large on a neighborhood, then the function must decrease significantly. This is the content of the following theorem; see [37, Basic Lemma, Chapter 1]) or [34, Lemma 2.5].
@@
Consider a lower semicontinuous function $f\colon\mathcal{X}\to\mathbf{R}\cup\{+\infty\}$ on a complete metric space $(\mathcal{X},\mathsf{d})$. Fix a point $x$ with $f(x)$ finite, and suppose that there are constants $\alpha<f(x)$ and $r,\kappa>0$ so that the implication holds:
@@
If in addition the inequality $f(x)-\alpha<\kappa r$ is valid, then the estimate holds:
@@
We will apply this theorem to the function $\mathcal{J}(\mu)^{1-q}$. The key step therefore is to compute the slope of $\mathcal{J}$. This is the content of the following lemma.
@@
Suppose that $F$ satisfies $\|DF(x)\|_{\rm op}\leq L(1+\|x\|_{2})$ for all $x\in\mathbf{R}^{d}$, where $L$ is some constant. Then, for any measure $\mu\in\mathcal{P}_{2}$, the estimate holds:
$|\nabla\mathcal{J}|(\mu)\geq\sup_{u\in E_{k}(\mu)}\sqrt{\operatorname{\mathbb{E}}_{\mu}\|DF(x)^{*}[uu^{\top}]\|^{2}_{2}}.$
@@
We begin by writing $\mathcal{J}(\mu)=(\lambda_{\min}\circ G)(\mu)$, where we define the map $G(\mu):=\operatorname{\mathbb{E}}_{\mu}F(x)$. Next, fix a measure $\mu\in\mathcal{P}_{2}$ and a matrix $U\in\partial\lambda_{\min}(G(\mu))$, and define the transport map $T(x)=x-DF(x)^{*}[U]$. Clearly, we may assume that $DF(x)^{*}[U]$ is not $\mu$-almost surely zero, since otherwise the theorem holds trivially for $U=uu^{\top}$. Observe that $I-T$ is square $\mu$-integrable since
@@
Define now the curve $\gamma\colon[0,1)\to\mathcal{P}_{2}$ by setting $\gamma(t)=(I+t(T-I))_{\#}\mu$. Note from (41), we have
@@
Next, from concavity of $\lambda_{\min}$ we deduce
@@
We would like to compute $\frac{d}{dt}\langle U,G\circ\gamma(t)\rangle$ by exchanging integration/differentiation in the expression:
@@
$|\langle U,DF(x+t(T(x)-x))[T(x)-x]\rangle|$ $\leq\|DF(x+t(T(x)-x))[T(x)-x]\|_{2}$
$\leq L(1+\|x+t(T(x)-x)\|_{2}\cdot\|T(x)-x\|_{2})$
$\leq L(1+\|x\|+t\|T(x)-x\|_{2})\|T(x)-x\|_{2}$
$\leq L\|T(x)-x\|_{2}+\frac{L}{2}\|x\|_{2}^{2}+\frac{L+2}{2}\|x-T(x)\|^{2}_{2}.$
Clearly, the right-side is $\mu$-integrable and therefore by the dominated convergence theorem, we may exchange integration and differentiation in (84) yielding:
@@
In particular, we deduce $(\mathcal{J}\circ\gamma)(t)<(\mathcal{J}\circ\gamma)(0)$ for all small $t>0$. Therefore, dividing (83) by $W_{2}(\gamma(t),\gamma(0))$ and taking the limit as $t\to 0$ yields
$|\nabla\mathcal{J}|(\mu)$ $\geq\lim_{t\to 0}\frac{\mathcal{J}(\gamma(0))-\mathcal{J}(\gamma(t))}{W_{2}(\gamma(0),\gamma(t))}$ (86)
$\geq\lim_{t\to 0}\frac{\mathcal{J}(\gamma(0))-\mathcal{J}(\gamma(t))}{t\sqrt{\operatorname{\mathbb{E}}_{\mu}\|x-T(x)\|^{2}_{2}}}$ (87)
$=\left\langle V,\lim_{t\to 0}\frac{(G\circ\gamma)(0)-(G\circ\gamma)(t)}{t\sqrt{\operatorname{\mathbb{E}}_{\mu}\|x-T(x)\|^{2}_{2}}}\right\rangle$ (88)
$=\sqrt{\operatorname{\mathbb{E}}_{\mu}\|T(x)-x\|^{2}_{2}}=\sqrt{\operatorname{\mathbb{E}}_{\mu}\|DF(x)^{*}[U]\|^{2}_{2}},$ (89)
where the estimate (87) follows from (82) and the estimate (89) follows from (85). ∎
Finally, combining the decrease principle (Theorem B.3) and the estimate on the slope of $\mathcal{J}$ (Lemma B.4) we arrive at the main result.
@@
Suppose that $F\colon\mathbf{R}^{d}\to\bm{S}^{d}_{+}$ satisfies $\|DF(x)\|_{\rm op}\leq L(1+\|x\|_{2})$ for all $x\in\mathbf{R}^{d}$, where $L$ is some constant. Fix a constant $c>0$, a radius $r>0$, and a power $q\in[0,1)$. Suppose that for all measures $\nu\in[0<\mathcal{J}\leq\mathcal{J}(\mu)]\cap\mathbb{B}_{2}(\mu;r)$, the estimate holds:
$\sup_{u\in E_{k}(\nu)}\operatorname{\mathbb{E}}_{\nu}\|DF(x)^{*}[uu^{\top}]\|^{2}_{2}\geq c\cdot\lambda_{\min}(\operatorname{\mathbb{E}}_{\nu}F(x))^{2q}.$
Then, the inequality holds:
$W_{2}(\mu,[\mathcal{J}=0])\leq\tfrac{1}{(1-q)\sqrt{c}}\cdot\mathcal{J}(\mu)^{1-q},$
as long as $r$ is large enough so that $\mathcal{J}(\mu)^{1-q}<(1-q)r\sqrt{c}$.
@@
It follows immediately from [66, Theorem 6.9] and continuity of $\lambda_{\min}(\cdot)$ that the function $\mathcal{J}$ is continuous. Define the function $\mathcal{G}(\nu):=\mathcal{J}(\nu)^{1-q}$ and note that the standard chain rule implies
$|\nabla\mathcal{G}|(\nu)=(1-q)\cdot\frac{|\nabla\mathcal{J}|(\nu)}{\mathcal{J}(\nu)^{q}}\geq(1-q)\sqrt{c},$

### v2

@@ B.2 Proof of Theorem 2
In this section, we prove Theorem 2, or rather a stronger version thereof. To this end, suppose there exists a $C^{1}$-smooth map $F\colon\mathbf{R}^{d}\to\bm{S}^{k}_{+}$ such that the Hessian $\nabla^{2}_{\mathcal{M}}f({\beta}_{\star})$ corresponding to any measure $\mu\in\mathcal{P}_{2}$ can be written as $\nabla^{2}_{\mathcal{M}}f({\beta}_{\star})={\operatorname{\mathbb{E}}}_{\mu}F(x)$. The differential of $F$ at $x$ will be denoted by $DF(x)\colon\mathbf{R}^{d}\to\bm{S}^{k}$, while the symbol $DF(x)^{*}\colon\bm{S}^{k}\to\mathbf{R}^{d}$ will denote the adjoint linear map of $DF(x)$. We assume that there exists a constant $L>0$ satisfying
$\|DF(x)\|_{\rm op}\leq L(1+\|x\|_{2})\qquad\forall x\in\mathbf{R}^{d}.$
By the mean value theorem, there exists a constant $C>0$ such that $\|F(x)\|_{F}\leq C(1+\|x\|_{2}^{2})$ for all $x\in\mathbf{R}^{d}$. Now define the function $\mathcal{J}\colon\mathcal{P}_{2}\to\mathbf{R}$ by setting
@@
It follows immediately from [70, Theorem 6.9] and the continuity of the function $\lambda_{\min}$ that $\mathcal{J}$ is continuous; in particular, the zero set $\mathcal{E}=[\mathcal{J}=0]$ is a closed subset of $\mathcal{P}_{2}$.
In order to simplify notation, for any matrix $A\in\bm{S}^{k}$ we let $E_{k}(A)$ denote the set of all unit eigenvectors of $A$ corresponding to the minimal eigenvalue $\lambda_{\min}(A)$. For any measure $\mu\in\mathcal{P}_{2}$, we define $E_{k}(\mu):=E_{k}({\operatorname{\mathbb{E}}}_{\mu}F(x))$. We will use the elementary facts that $\lambda_{\min}$ is a Lipschitz continuous concave function on $\bm{S}^{k}$ and its supdifferential at any $A\in\bm{S}^{k}$ is the set
$\partial\lambda_{\min}(A)=\operatorname{conv}\{uu^{\top}\mid u\in E_{k}(A)\};$
see, for example, [9, Corollary 5.2.3, Corollary 5.2.4 (iii)].
The proof of Theorem 2 will be divided into two parts, corresponding to the two inequalities in (15). We begin by establishing the first inequality ${\rm REG}(\mu)^{q_{2}-1}\lesssim{\rm RSE}(\mu)$; this is the content of the following theorem. The proof amounts to applying the fundamental theorem of calculus to a power of the function $\mathcal{J}$ along a geodesic $\mu_{t}$ joining a measure $\mu\in\mathcal{P}_{2}\setminus\mathcal{E}$ to a nearby measure in $\mathcal{E}$.
@@
Fix a measure $\mu\in\mathcal{P}_{2}\setminus\mathcal{E}$, constants $c,\varepsilon>0$, and a power $q\in[0,1)$. Suppose that for all measures $\nu\in\mathcal{P}_{2}\setminus\mathcal{E}$ satisfying $W_{2}(\mu,\nu)<W_{2}(\mu,\mathcal{E})+\varepsilon$, the estimate holds:
$\min_{u\in E_{k}(\nu)}{\operatorname{\mathbb{E}}}_{\nu}\|DF(x)^{*}[uu^{\top}]\|^{2}_{2}\leq c\cdot\mathcal{J}(\nu)^{2q}.$
Then
$\mathcal{J}(\mu)^{1-q}\leq\sqrt{c}\,(1-q)\cdot W_{2}(\mu,\mathcal{E}).$
@@
Given any $\delta\in(0,\varepsilon]$, select a measure $\bar{\nu}\in\mathcal{E}$ satisfying $W_{2}(\mu,\bar{\nu})<W_{2}(\mu,\mathcal{E})+\delta$, and let $\pi\in\Pi(\bar{\nu},\mu)$ be an optimal transport plan from $\bar{\nu}$ to $\mu$. Define the functions $\rho_{t}(x,y)=(1-t)x+ty$ for all $t\in[0,1]$ and $x,y\in\mathbf{R}^{d}$. Then the curve $\mu_{t}=(\rho_{t})_{\#}\pi$ is a constant speed geodesic in $\mathcal{P}_{2}$ from $\bar{\nu}$ to $\mu$ [62, Theorem 5.27]:
$W_{2}(\mu_{t},\mu_{s})=|t-s|\cdot W_{2}(\mu,\bar{\nu})\qquad\forall t,s\in[0,1].$
Now define the curve $\gamma\colon[0,1]\to\bm{S}_{+}^{k}$ by setting $\gamma(t)={\operatorname{\mathbb{E}}}_{\mu_{t}}F(x)$. We would like to compute $\dot{\gamma}(t)$ by interchanging differentiation and integration in the expression
$\dot{\gamma}(t)=\frac{d}{dt}\,{\operatorname{\mathbb{E}}}_{\pi}F((1-t)x+ty).$ (78)
To this end, we bound the derivative of the integrand uniformly in $t$:
$\|DF((1-t)x+ty)[y-x]\|_{F}\leq L(1+\|x\|_{2}+\|y\|_{2})\|y-x\|_{2}.$
Applying Hölder’s inequality, we see that this upper bound is $\pi$-integrable. Therefore, interchanging differentiation and integration in (78) yields the expression $\dot{\gamma}(t)={\operatorname{\mathbb{E}}}_{\pi}DF((1-t)x+ty)[y-x]$.
It is clear that $\gamma$ is absolutely continuous and therefore, using the supdifferential chain rule for concave functions [11, Lemma 3.3, p. 73], we deduce that for almost every $t\in(0,1)$ we have
@@
Define $t_{0}:=\max\{t\in[0,1]\mid\mathcal{J}(\mu_{t})=0\}$ and observe $t_{0}\in[0,1)$ and, for all $t\in(t_{0},1]$, we have $\mu_{t}\in\mathcal{P}_{2}\setminus\mathcal{E}$ and $W_{2}(\mu,\mu_{t})=(1-t)\cdot W_{2}(\mu,\bar{\nu})<W_{2}(\mu,\mathcal{E})+\varepsilon.$ Thus, by assumption, for each $t\in(t_{0},1)$, we may choose $U_{t}\in\partial\lambda_{\min}(\gamma(t))$ satisfying ${\operatorname{\mathbb{E}}}_{\mu_{t}}\|DF(x)^{*}[U_{t}]\|^{2}_{2}\leq c\cdot\mathcal{J}(\mu_{t})^{2q}$. Continuing with (79), we successively compute
$\frac{d}{dt}\mathcal{J}(\mu_{t})$ $={\operatorname{\mathbb{E}}}_{\pi}\langle DF((1-t)x+ty)^{*}[U_{t}],y-x\rangle$ (80)
$\leq{\operatorname{\mathbb{E}}}_{\pi}\big[\|DF((1-t)x+ty)^{*}[U_{t}]\|_{2}\cdot\|y-x\|_{2}\big]$ (81)
$\leq\sqrt{{\operatorname{\mathbb{E}}}_{\pi}\|DF((1-t)x+ty)^{*}[U_{t}]\|^{2}_{2}}\cdot\sqrt{{\operatorname{\mathbb{E}}}_{\pi}\|y-x\|_{2}^{2}}$ (82)
$\leq\sqrt{c}\cdot W_{2}(\mu,\bar{\nu})\cdot\mathcal{J}(\mu_{t})^{q}$ (83)
for a.e. $t\in(t_{0},1)$, where (82) follows from Hölder’s inequality. Raising $\mathcal{J}(\mu_{t})$ to power 1q1-q, we deduce
$\frac{d}{dt}\mathcal{J}(\mu_{t})^{1-q}\leq\sqrt{c}\,(1-q)\cdot W_{2}(\mu,\bar{\nu})\qquad\textrm{for a.e. }t\in(t_{0},1).$
Integrating both sides from $t=t_{0}$ to $t=1$ and leveraging absolute continuity, we conclude
$\mathcal{J}(\mu)^{1-q}-\mathcal{J}(\mu_{t_{0}})^{1-q}\leq(1-t_{0})\cdot\sqrt{c}\,(1-q)\cdot W_{2}(\mu,\bar{\nu}).$ (84)
Taking into account the equality $\mathcal{J}(\mu_{t_{0}})=0$ and the estimate $W_{2}(\mu,\bar{\nu})<W_{2}(\mu,\mathcal{E})+\delta$, we obtain
$\mathcal{J}(\mu)^{1-q}<\sqrt{c}\,(1-q)\cdot\big(W_{2}(\mu,\mathcal{E})+\delta\big).$
This last inequality holds for all $\delta\in(0,\varepsilon]$, demonstrating $\mathcal{J}(\mu)^{1-q}\leq\sqrt{c}\,(1-q)\cdot W_{2}(\mu,\mathcal{E})$. ∎
@@
$|\nabla f|(x):=\limsup_{x^{\prime}\to x}\frac{(f(x)-f(x^{\prime}))_{+}}{\mathsf{d}(x,x^{\prime})}.$
Importantly, if the slope is large on a neighborhood, then the function must decrease significantly. This is the content of the following theorem; see [38, Basic Lemma, Chapter 1] or [33, Lemma 2.5].
@@
Consider a lower semicontinuous function $f\colon\mathcal{X}\to\mathbf{R}\cup\{+\infty\}$ on a complete metric space $(\mathcal{X},\mathsf{d})$. Fix a point $x$ with $f(x)$ finite, and suppose that there are constants $\alpha<f(x)$ and $r,\kappa>0$ such that the implication holds:
@@
If the inequality $f(x)-\alpha<\kappa r$ is valid, then the estimate holds:
@@
We will apply this theorem to a power of the function $\mathcal{J}$. The key step therefore is to estimate the slope of $\mathcal{J}$ (with respect to the metric $W_{2}$ on $\mathcal{P}_{2}$). This is the content of the following lemma.
@@
For any measure $\mu\in\mathcal{P}_{2}$, the estimate holds:
$|\nabla\mathcal{J}|(\mu)\geq\max_{u\in E_{k}(\mu)}\sqrt{{\operatorname{\mathbb{E}}}_{\mu}\|DF(x)^{*}[uu^{\top}]\|^{2}_{2}}.$
@@
We begin by writing $\mathcal{J}(\mu)=(\lambda_{\min}\circ G)(\mu)$, where we define the map $G(\mu):={\operatorname{\mathbb{E}}}_{\mu}F(x)$. Next, fix a measure $\mu\in\mathcal{P}_{2}$ and a matrix $U\in\partial\lambda_{\min}(G(\mu))$, and define the transport map $T(x):=x-DF(x)^{*}[U]$. Observe that ITI-T is square $\mu$-integrable since
@@
Define now the curve $\gamma\colon[0,1]\to\mathcal{P}_{2}$ by setting $\gamma(t)=(I+t(T-I))_{\#}\mu$. Note that from (44), we have
@@
Next, from the concavity of $\lambda_{\min}$ we deduce
@@
We would like to compute $\frac{d}{dt}\langle U,G\circ\gamma(t)\rangle$ by interchanging differentiation and integration in the expression
@@
$|\langle U,DF(x+t(T(x)-x))[T(x)-x]\rangle|$ $\leq\|U\|_{F}\cdot\|DF(x+t(T(x)-x))[T(x)-x]\|_{F}$
$\leq L(1+\|x+t(T(x)-x)\|_{2})\cdot\|T(x)-x\|_{2}$
$\leq L(1+\|x\|_{2}+t\|T(x)-x\|_{2})\cdot\|T(x)-x\|_{2}$
$\leq L\Big(\|T(x)-x\|_{2}+\tfrac{1}{2}\|x\|_{2}^{2}+\tfrac{3}{2}\|x-T(x)\|^{2}_{2}\Big).$
Clearly, the right-side is $\mu$-integrable and therefore by the dominated convergence theorem, we may interchange differentiation and integration in (87), yielding:
@@
In particular, we deduce $(\mathcal{J}\circ\gamma)(t)<(\mathcal{J}\circ\gamma)(0)$ for all small $t>0$. Therefore, dividing (86) by $W_{2}(\gamma(t),\gamma(0))$ and taking the limit as $t\downarrow 0$ yields
$|\nabla\mathcal{J}|(\mu)$ $\geq\lim_{t\downarrow 0}\frac{\mathcal{J}(\gamma(0))-\mathcal{J}(\gamma(t))}{W_{2}(\gamma(0),\gamma(t))}$ (89)
$\geq\lim_{t\downarrow 0}\frac{\mathcal{J}(\gamma(0))-\mathcal{J}(\gamma(t))}{t\sqrt{{\operatorname{\mathbb{E}}}_{\mu}\|x-T(x)\|^{2}_{2}}}$ (90)
$\geq\left\langle U,\lim_{t\downarrow 0}\frac{(G\circ\gamma)(0)-(G\circ\gamma)(t)}{t\sqrt{{\operatorname{\mathbb{E}}}_{\mu}\|x-T(x)\|^{2}_{2}}}\right\rangle$ (91)
$=\sqrt{{\operatorname{\mathbb{E}}}_{\mu}\|T(x)-x\|^{2}_{2}}$ (92)
$=\sqrt{{\operatorname{\mathbb{E}}}_{\mu}\|DF(x)^{*}[U]\|^{2}_{2}},$ (93)
where the estimate (90) follows from (85) and the equality (92) follows from (88). ∎
Finally, combining the decrease principle (Theorem B.3) and the estimate on the slope of $\mathcal{J}$ (Lemma B.4), we arrive at the main result.
@@
Fix a measure $\mu\in\mathcal{P}_{2}\setminus\mathcal{E}$, a constant $c>0$, a radius $r>0$, and a power $q\in[0,1)$. Suppose that for all measures $\nu\in[0<\mathcal{J}\leq\mathcal{J}(\mu)]\cap\overline{\mathbb{B}}_{2}(\mu;r)$, the estimate holds:
$\max_{u\in E_{k}(\nu)}{\operatorname{\mathbb{E}}}_{\nu}\|DF(x)^{*}[uu^{\top}]\|^{2}_{2}\geq c\cdot\lambda_{\min}({\operatorname{\mathbb{E}}}_{\nu}F(x))^{2q}.$
Then
$W_{2}(\mu,[\mathcal{J}=0])\leq\tfrac{1}{(1-q)\sqrt{c}}\cdot\mathcal{J}(\mu)^{1-q}$
provided $r$ is large enough so that $\mathcal{J}(\mu)^{1-q}<r(1-q)\sqrt{c}$.
@@
Define the function $\mathcal{G}(\nu):=\mathcal{J}(\nu)^{1-q}$ and note that the slope chain rule and Lemma B.4 imply
$|\nabla\mathcal{G}|(\nu)=(1-q)\cdot\frac{|\nabla\mathcal{J}|(\nu)}{\mathcal{J}(\nu)^{q}}\geq(1-q)\sqrt{c}$